For this non-homogeneous diff eq. what is the form of the particular solution that you would guess in the method of undetermined coefficients.

y + 2y + 2y = 4tet + 3te-t cos(t)
Answer #1

q+zyltzy =utet + 3t et cast Auxillary earstion : +20 +2=0 =, D = -1 ti Porti Cila solution up = CAٲ+ Bt tc) et + cott Etje